1. In bench pressing 100 kg, a weight lifter applies a 1040 N force to the bottom of the bar.  
1
Expert's answer
2021-11-08T09:12:06-0500

Fnet=maFmg=maa=Fmgm=10401009.811000.6 (m/s2)F_{net}=ma\to F-mg=ma\to a=\frac{F-mg}{m}=\frac{1040-100\cdot9.81}{100}\approx0.6 \ (m/s^2) . Answer